What is the ordered pair of the equations y=1/5x+3/5 y=4/7xa-2
Answer by fractalier(6550)

Please repost this one...not sure what the "a" is doing there...if it is a typo, then for
y=1/5x+3/5
y=4/7x-2
we have two expressions both equal to y...set them equal and start solving for x:
1/5x + 3/5 = 4/7x - 2